Way Corporation disposed of the following tangible personal property assets in the current year. Date Date Original Asset Acquired Sold Convention Basis Furniture (7-year) 5/12/13 7/15/19 HY $ 55,000 Machinery (7-year) 3/23/14 3/15/19 MQ 72,000 Delivery truck* (5-year) 9/17/16 3/13/19 HY 20,000 Machinery (7-year) 10/11/17 8/11/19 MQ 250,000 Computer (5-year) 2/11/19 12/15/19 HY 80,000 *Used 100 percent for business. Assume that the delivery truck is not a luxury auto. Calculate Way Corporation’s 2019 depreciation expense (Way has not used §179 expense and bonus depreciation in past years and elects not to do so this year as well). Be sure to show your work if you want any partial credit.
Answer #1

in the Answer the tangible personal property assets Current year Furniture (1-year) Dale acquered osta 2/20142 Date sold offMachinery (7-year) Sabe acquiel 03 ) 23/2015 Date sold 03/15/2018 Quarter - 1st Porion of the year = 12.50% original Baris =-) 50.001 20000 100.000 Rate = 19.20% = 100.000X19.20) . = 1920.00 Dep Esip. MachineryL7-year) Date acquired = 10/11/2017 Dal27.55% X 156.250 Sep fno 243046875 2 Computer Ls-year) Dale acqueried = 10/1/2018 Dale sold - 12/15/2018 . Quater : = n/a Per